# thanks: Present address

### v4

@@ IV Results
Figure 5 plots of $\tilde{F}_{\mathrm{max}}$ against $U_{0}$ for various packing fractions, $\phi=0.52,0.53,0.54$, and $0.56$. The blue solid circles indicate experimental estimations of $\tilde{F}_{\mathrm{max}}$, and the black solid lines are the numerical solutions of Eq. (4), where the dimensionless effective viscosity $\eta$ is treated as a fitting parameter with its error bar. The gray areas around the solid line indicate the possible solution within the error of estimated $\eta$. It is readily observed that the theoretical curvesclosely overlap with the experimental data points including their error bars. This agreement is particularly pronounced at higher packing fractions ($\phi=0.54$ and $0.56$), although the theoretical predictions successfully capture the overall trend of the experimental data even at lower packing fractions ($\phi=0.52$ and $0.53$). We also confirm the asymptotic behavior predicted from the floating model as $\tilde{F}_{\mathrm{max}}\sim U_{0}^{3/2}$ for $U_{0}\gg 1$ in Eq. 6 [11, 12]. The solid lines in Fig. 5 indicate the power exponents of the asymptotic solution, which agrees with the slope from the experimental data points. These results indicate that the experimental data across the entire range from $\phi=0.52$ to $\phi=0.56$ are well described by the solution of the floating model given in Eq. (4). Unfortunately, we cannot get definite results for $t_{\mathrm{max}}$, where $t_{\mathrm{max}}$ weakly depends on $U_{0}$. This tendency was reported even in the simulation [11]. We need further investigation on $t_{\mathrm{max}}$.
@@
Figure 6 exhibits how the dimensionless effective viscosity $\eta$ depends on $\phi$. The figure indicates that $\eta$ weakly depends on $\phi$ in the range $\phi=0.52-0.56$, where $\eta$ slightly increases with $\phi$. This result is unexpected because the viscosity in usual densely packed suspensions increases rapidly as the density approaches the jamming point or the random closed packing value. We will discuss this insensitivity of $\eta$ to $\phi$ in Sec. V.
@@ V Discussion
As presented in the previous section, we found that the floating model provides a good description of our experimental results. In particular, the behavior of the maximum drag force, $\tilde{F}_{\mathrm{max}}$, as a function of the impact speed $U_{0}$ shows reasonable agreement with the theoretical curves. These results validate several simplifications and assumptions made in the floating model, namely: the approximation of buoyancy forces (see Appendix B), the neglect of elastic forces, and the use of Stokes drag when focusing on the early stage of the dynamics. It is important to note that the elastic force becomes significant during the late stage of the dynamics, although its contribution remains much smaller than that of the viscous drag force during the early stage of impact [12].
@@
To evaluate potential sidewall effects due to the quasi-2D container used in the experiment, we also examine a cubic container (Appendix A). The experimental results for $F_{\mathrm{max}}$ versus $U_{0}$ in the cubic container (see Fig.7) are almost indistinguishable from those in Fig.5. This confirms the robustness of the floating model, regardless of the container shape.
Although several models have been proposed to describe impact processes in suspensions [4, 8, 9, 10, 11, 12], most of these models include an elastic term to capture the rebound behavior observed in the late stage of the impact. In contrast, the floating model is particularly well-suited for describing the early-stage dynamics, where viscous effects dominate. Therefore, we conclude that the floating model accurately captures the early-stage dynamics in dense suspensions of millimeter-sized particles.
@@
As shown in Fig. 6, the effective viscosity $\eta$ exhibits only weak dependence on the packing fraction $\phi$. This insensitivity can be attributed to two main factors. First, the actual packing fractions are higher than the nominal ones due to imperfect density matching; for instance, $\varphi=0.40$, 0.48, 0.51, and 0.56 correspond to $\phi=0.52$, 0.53, 0.54, and 0.56, respectively. As a result, our experiments cover only a narrow range of $\phi$. Second, empirical relations suggest that the increase in effective viscosity is not substantial within this narrow range, especially under the assumption that the viscosity diverges at the random close packing fraction $\phi_{\mathrm{rcp}}\approx 0.64$. For example, if we apply the sedimentation theory used for $\phi\leq 0.50$ [23], the ratio of effective viscosity between $\phi=0.56$ and $\phi=0.52$ is about 1.3. Even using empirical relations for denser, sheared non-Brownian suspensions [24], the ratio would be about 2.0. These estimates suggest that the increase in effective viscosity over our observed range is not significant, assuming divergence at $\phi_{\mathrm{rcp}}$, though a detailed quantitative justification remains an open question.
@@ VI Conclusions
In summary, this study demonstrates the applicability of the floating model [9, 11, 12] to suspensions containing millimeter-sized particles, particularly in capturing the relationship between the maximum drag force, $F_{\mathrm{max}}$, and the impact speed, $u_{0}$. Our experimental results confirm the scaling law $F_{\mathrm{max}}\sim u_{0}^{3/2}$ for $u_{0}\gg\sqrt{ga_{\rm I}}$, in agreement with theoretical predictions.
These findings support the validity of previous lattice Boltzmann method (LBM) simulations [11, 12] conducted on sparse suspensions, despite notable differences in the size ratio between the projectile and the suspended particles. Surprisingly, our results suggest that Stokes flow may effectively describe fluid-particle interactions even for large particles, under certain conditions.
Looking ahead, a deeper understanding of why the floating model applies to suspensions of millimeter-sized particles remains an important direction for future research. Although this study primarily focuses on early-stage dynamics, future work should explore the role of elastic forces in the late stages of impact, so-called the IIH as indicated by previous experimental [8] and theoretical [11, 12] studies.
@@ Appendix A The dynamical impact of experiments on the cubic cell
Figure 7: The plots between the dimensionless maximum force $\tilde{F}_{\mathrm{max}}:=F_{\mathrm{max}}/m_{\rm I}g$ and impact-speed $U_{0}:=u_{0}/\sqrt{a_{\rm I}g}$ for the cubic cell in the packing fraction $\phi=0.56$. The red solid circles indicate the experimental results and the black solid lines are obtained by the numerical solutions of the floating model Eq. (4), where $\eta=12.0\pm 1.2$. The gray area around the solid line indicates the possible solutions within the errors of the estimated $\eta$. The black solid lines in large $U_{0}$ are the guide lines $\tilde{F}_{\mathrm{max}}\propto U_{0}^{3/2}$ as expected from Eq. (6).
@@ Appendix B The expression of the buoyancy force
This expression is valid for $-2a_{\rm I}\leq z\leq 0$.
@@
$=$ $-m_{\rm I}g\left\{1-\frac{\rho_{\rm f}}{4\ \rho_{\rm I}}\left[3\left(\frac{z}{a_{\rm I}}\right)^{2}-\left(\frac{|z|}{a_{\rm I}}\right)^{3}\right]\right\}$
@@ Appendix C Exact solution of floating model
[…]

### v5

@@ IV Results
Figure 6 plots of $\tilde{F}_{\mathrm{max}}$ against $U_{0}$ for various packing fractions, $\Phi=0.52,0.53,0.54$, and $0.56$. The blue solid circles indicate experimental estimations of $\tilde{F}_{\mathrm{max}}$, and the dark-gray solid lines are the numerical solutions of Eq. (4), where the dimensionless effective viscosity $\eta$ is treated as a fitting parameter with its error bar. The gray areas around the solid line indicate the possible solution within the error of the estimated $\eta$. It is readily observed that the theoretical curves closely overlap with the experimental data points, including their error bars. This agreement is particularly pronounced at higher packing fractions ($\Phi=0.54$ and $0.56$), although the theoretical predictions successfully capture the overall trend of the experimental data even at lower packing fractions ($\Phi=0.52$ and $0.53$). We also confirm the asymptotic behavior predicted from the floating model as $\tilde{F}_{\mathrm{max}}\sim U_{0}^{3/2}$ for $\eta U_{0}\gg 1$ in Eq. (6) [43, 44]. The black solid line segments in Fig. 6 indicate the power exponents of the asymptotic solution, which agree with the slope from the experimental data points. These results indicate that the experimental data across the entire range from $\Phi=0.52$ to $\Phi=0.56$ are well described by the solution of the floating model given in Eq. (4).
@@
Figure 7 exhibits how the dimensionless effective viscosity $\eta$ depends on $\Phi$. The figure indicates that $\eta$ weakly depends on $\Phi$ in the range $\Phi=0.52-0.56$, where $\eta$ slightly increases with $\Phi$. This result is unexpected because the viscosity in usual densely packed suspensions increases rapidly as the density approaches the jamming point or the random closed packing value. Since the fraction in the dynamical jammed region determines the effective viscosity, a lower $\Phi$ has a larger effective packing fraction. We will discuss this insensitivity of $\eta$ to $\Phi$ in Sec. V.
@@ V Discussion
As presented in the previous section, we found that the floating model provides a good description of our experimental results. In particular, the behavior of the maximum drag force, $\tilde{F}_{\mathrm{max}}$, as a function of the impact speed $U_{0}$ shows reasonable agreement with the theoretical curves. These results validate several simplifications and assumptions made in the floating model, namely, the approximation of buoyancy forces (see Appendix B), the neglect of elastic forces, and the use of Stokes drag when focusing on the early stage of the dynamics. It is important to note that the elastic force becomes significant during the late stage of the dynamics, although its contribution remains much smaller than that of the viscous drag force during the early stage of impact [41, 43, 44].
Although several models have been proposed to describe impact processes in suspensions [35, 40, 41, 42, 43, 44], most of these models include an elastic term to capture the rebound behavior observed in the late stage of the impact. In other words, the elastic term strongly depends on the contact force between suspended particles, such as interparticle friction [42]. In contrast, the floating model is particularly well-suited for describing the early-stage dynamics, where viscous effects dominate, and is insensitive to the contact force between particles. Therefore, we conclude that the floating model accurately captures the early-stage dynamics in dense suspensions of millimeter-sized particles.
@@
Figure 8: The plots between $\tau_{\mathrm{max}}$ and $U_{0}$ for various $\Phi=0.52,0,53,0.54$, and 0.56, where the vertical lines are error bars in the estimation of $\tau_{\mathrm{max}}$.
Unfortunately, our results for $\tau_{\mathrm{max}}$ is inconsistent with the prediction of the floating model [43], where $\tau_{\mathrm{max}}$ weakly depends on $U_{0}$ (see Fig. 8), while the floating model predicts $\tau_{\mathrm{max}}\sim U_{0}^{-1/2}$ for large $U_{0}$. However, the previous experiment and simulation also do not get perfect agreement with the theoretical prediction [41, 43]. This is partially because the floating model contains only one fitting parameter $\eta$ and has a quantitative deviation of the estimated $\tau_{\mathrm{max}}$ between the experiment and the theoretical one, as shown in Fig. 5. More importantly, the onset of the drag force in the experiments may be governed by the mechanisms which is not involved in the floating model.
Figure 9: Scaling plots for normalized drag force $F/F_{\mathrm{max}}$ versus the dimensionless time $(t-t_{\mathrm{max}})/(t_{\mathrm{max}}-t_{\mathrm{half}})$ for various impact speeds at $\Phi=0.56$, where symbols show experimental observations and the solid lines are the solutions of the floating model for various $U_{0}$ with $\eta=9.06$ at the corresponding impact speeds.
To clarify the time-dependent properties of the impact process, we introduce the half time, $t_{\mathrm{half}}$ to reach $F_{\mathrm{max}}/2$ where $\tau_{\mathrm{half}}=t_{\mathrm{half}}\sqrt{g/a_{\mathrm{I}}}$. [^2] Similar to Fig. 8, the behavior of $\tau_{\mathrm{half}}$ versus $U_{0}$ cannot be described by the floating model. Figure 9 presents scaling plots of $F/F_{\mathrm{max}}$ versus $(t-t_{\mathrm{max}})/(t_{\mathrm{max}}-t_{\mathrm{half}})$ for various initial speeds $U_{0}$. Experimental data for $(t-t_{\mathrm{max}})/(t_{\mathrm{max}}-t_{\mathrm{half}})<0$ and the solution of the floating model exhibit reasonable data collapse to support the scaling hypothesis near the peak. Note that the elastic force plays an important role for $(t-t_{\mathrm{max}})/(t_{\mathrm{max}}-t_{\mathrm{half}})>0$. It should be noted that the scaling hypothesis for the experimental data is no longer valid near the point of impact of the projectile on the liquid, although the floating model satisfies the scaling for all time regions. This indicates that another mechanism is necessary to explain Fig. 8.
@@
To evaluate sidewall effects due to the quasi-2D container used in the experiment, we also examine a cubic container (Appendix A). The experimental results for $F_{\mathrm{max}}$ versus $U_{0}$ in the cubic container (see Fig.10) are almost indistinguishable from those in Fig.6. This confirms the robustness of the floating model, regardless of the container shape.
We also examine several other situations in which the diameter of the projectile is set to be $D_{\mathrm{I}}=8$ mm instead of $D_{\mathrm{I}}=12$ mm as in the main text, and we have also examined a viscous liquid with 9 times larger viscosity using a water-glycerol suspension (see Appendix D). Interestingly, we have confirmed that our results reported in the main text are unchanged even when we use smaller projectiles and/or viscous solvents. These results suggest that our results in this paper are robust for relatively general situations.
@@
As shown in Fig. 7, the effective viscosity $\eta$ exhibits only weak dependence on the packing fraction $\Phi$. This insensitivity can be attributed to two main factors. First, the actual packing fractions are higher than the nominal ones due to imperfect density matching; for instance, $\varphi=0.40$, 0.48, 0.51, and 0.56 correspond to $\Phi=0.52$, 0.53, 0.54, and 0.56, respectively. As a result, our experiments cover only a narrow range of $\Phi$. Second, empirical relations suggest that the increase in effective viscosity is not substantial within this narrow range, especially under the assumption that the viscosity diverges at the random close packing fraction $\Phi_{\mathrm{rcp}}\approx 0.64$. For example, using empirical relations for denser, sheared non-Brownian suspensions [13, 15], the ratio is at most 2.0, although our system is not sheared. If we apply the sedimentation theory[12] used for $\Phi\leq 0.50$, the ratio of effective viscosity between $\Phi=0.56$ and $\Phi=0.52$ is about 1.3, though this theory cannot be used for $\Phi>0.50$. According to Appendix E, the extended sedimentation theory of Ref. [12] for denser regions ($\Phi>0.50$) suggests that the ratio is merely 1.07. Since the impact process is not sheared but is related to sedimentation, we should adopt the extended sedimentation theory to understand $\Phi$ dependence of $\eta$, which is insensitive to $\Phi$ in our observed region $0.52\leq\Phi\leq 0.56$, These estimates suggest that the increase in effective viscosity over our observed range is not significant, assuming divergence at $\Phi_{\mathrm{rcp}}$, though a detailed quantitative justification remains an open question.
@@ VI Conclusions
We have experimentally demonstrated that the floating model, originally developed for micrometer-sized particle suspensions, remains valid in describing the early-stage impact dynamics (for $t_{\mathrm{half}}<t<t_{\mathrm{max}}$) of millimeter-sized particle suspensions. The observed scaling law $F_{\mathrm{max}}\sim u_{0}^{3/2}$ for $\eta U_{0}\gg 1$ aligns with theoretical predictions, confirming the dominance of viscous drag in the initial phase of impact. Our findings extend the applicability of the viscous-based floating model and suggest that Stokes flow assumptions may hold under specific conditions, even for larger particles. This work bridges the gap between microscale simulations and macroscale experiments, offering new insights into the mechanics of dense suspensions.
Looking ahead, a deeper understanding of why the floating model applies to suspensions of millimeter-sized particles remains an important direction for future research. Although this study primarily focuses on early-stage dynamics, future work should explore the role of elastic forces in the late stages of impact, so-called the IIH, as indicated by previous experimental [40] and theoretical [43, 44] studies. The floating model fails to capture the elastic contribution at the instant of impact between the projectile and the particles in the suspension. Elucidating these dynamics remains an important direction for future work. Also, using large-sized particles, we may visualize the force propagation in force chains in the suspended particles in the forthcoming study.
@@ Appendix A The dynamical impact of experiments on the cubic cell
Figure 10: The plots between the dimensionless maximum force $\tilde{F}_{\mathrm{max}}:=F_{\mathrm{max}}/m_{\rm I}g$ and impact-speed $U_{0}:=u_{0}/\sqrt{a_{\rm I}g}$ for the cubic cell in the packing fraction $\Phi=0.56$. The red solid circles indicate the experimental results and the black solid lines are obtained by the numerical solutions of the floating model Eq. (4), where $\eta=10.09\pm 1.90$. The gray area around the solid line indicates the possible solutions within the errors of the estimated $\eta$. The black solid line segments in large $U_{0}$ are the guide lines $\tilde{F}_{\mathrm{max}}\propto U_{0}^{3/2}$ as expected from Eq. (6).
@@ Appendix B The expression of the buoyancy force
This expression is valid for $2a_{\rm I}\leq z\leq 0$.
@@
$=$ $-m_{\rm I}g\left\{1-\frac{\rho_{\rm f}}{4\ \rho_{\rm I}}\left[3\left(\frac{z}{a_{\rm I}}\right)^{2}-\left(\frac{|z|}{a_{\rm I}}\right)^{3}\right]\right\},$
@@ Appendix C Exact solution of floating model
where $\Xi:=\left(\frac{\eta\left(1-\xi\right)}{2}\right)^{1/3}\left(\tau+\frac{U_{0}}{1-\xi}\right)$, $\Lambda:=\left(\frac{\eta}{2\left(1-\xi\right)^{2}}\right)^{1/3}U_{0}$. ${\rm Ai}\left(x\right)$ and ${\rm Bi}\left(x\right)$ are Airy functions as ${\rm Ai}\left(x\right):=\frac{1}{\pi}\int_{0}^{\infty}\cos\left(\frac{t^{3}}{3}+xt\right)dt$, and ${\rm Bi}\left(x\right):=\frac{1}{\pi}\int_{0}^{\infty}\left[e^{-\frac{t^{3}}{3}+xt}+\sin\left(\frac{t^{3}}{3}+xt\right)\right]dt$, ${\rm Ai}^{\prime}(x)$ and ${\rm Bi}^{\prime}(x)$ are their derivatives with respect to $x$, respectively. Note that $F_{\mathrm{max}}$ cannot be obtained exactly, although the asymptotic expression for $U_{0}\gg 1$ can be evaluated [43].
Appendix D Comparison between various solution viscosities and projectile sizes.
We compare the plots of $F_{\mathrm{max}}$ versus $U_{0}$ using a water-glycerol suspension, and a smaller projectile in Fig. 12. The suspension solution is prepared by mixing with water and glycerol (water:glycerol=1:1 in volume fraction), whose viscosity is increased to be about 9 times larger than water, $\eta_{0}=8.8~{\rm mPa~s}$. Then the glycerol/water solution is mixed with particles with the packing fraction $\Phi=0.56$. As the density of the glycerol/water solution becomes $1140~{\rm kg/m^{3}}$, the density matching is impossible to achieve in this case. In a separate experiment, a smaller projectile with a diameter $D=8~\mathrm{mm}$ is used in a water-based suspension with the same packing fraction ($\Phi=0.56$). Remarkably, all the data points collapse onto a single curve, exhibiting a power-law scaling with an exponent of $3/2$ at higher impact speeds.
The effective viscosity is estimated as $\eta=8.99\pm 1.30$ for the suspension of water/glycerol, $\eta=6.88\pm 1.23$ for a smaller projectile $D=8~{\rm mm}$. In comparison with water suspension, $\eta=9.06\pm 1.45$, the data plots are seemingly overlapping. Thus, the floating model can be used for suspensions with smaller projectiles and/or viscous solvents.
Figure 12: A comparison of the $\tilde{F}_{\mathrm{max}}$$U_{0}$ plots for the water suspension with $D=12~\mathrm{mm}$, the water/glycerol suspension with $D=12~\mathrm{mm}$, and water suspension with $D=8~\mathrm{mm}$. All the packing fraction is $\Phi=0.56$.
Appendix E Extended sedimentation theory
In this appendix, we attempt to extend the sedimentation theory for $\Phi<0.49$ [12] to the denser situation, $\Phi>0.49$, using an empirical argument. The sedimentation rate for one particle sedimentation rate $U_{0}^{\mathrm{sed}}$, in a fluid is expressed as
$U^{\mathrm{sed}}_{0}:=\frac{2}{9\eta_{0}}\Delta\rho ga_{\mathrm{I}}^{2},$ (10)
where $\eta_{0}$ is the viscosity of the solvent, $\Delta\rho:=\rho_{\mathrm{I}}-\rho_{\mathrm{f}}$ and $a_{\mathrm{I}}:=d_{\mathrm{I}}/2$. If we are interested in the sedimentation rate in random suspensions with the packing fraction $\Phi$, this can be written as
$\frac{U^{\mathrm{sed}}(\Phi)}{U_{0}^{\mathrm{sed}}}=\frac{\eta_{0}}{\eta_{\mathrm{eff}}}\propto\eta^{-1}.$ (11)
Therefore, to estimate the viscosity in the dense random suspensions is to estimate $U_{0}^{\mathrm{sed}}(\Phi)$. It is known that the sedimentation rate is the evaluation of the ensemble average of the mobility matrix $M$, i.e.
$\langle M\rangle=\frac{U^{\mathrm{sed}}(\Phi)}{U_{0}^{\mathrm{sed}}}\propto\eta^{-1}.$ (12)
The sedimentation theory in Ref. [12] consists of two parts: one is from the long-range Rotne-Prager part, and the other is the lubrication part. Both contributions contain parts proportional to the density and the convolution integrals containing the pair correlation function. The lubrication part is unaffected by the density. Thus, the averaged lubrication matrix over random suspensions $\langle\mathsf{R}^{\mathrm{lub}}\rangle$ can be evaluated as $\langle\mathsf{R}^{\mathrm{lub}}\rangle\approx 1.492\Phi$[12]
On the other hand, Brady and Durlofsky [11] evaluated the long-range part, the random average of the mobility matrix $\langle M^{\infty}\rangle$ as
$\langle M^{\infty}\rangle=1-5\Phi-\frac{\Phi^{2}}{5}+3\Phi\int_{2}^{\infty}dr^{\prime}r^{\prime}\{g_{2}(r^{\prime})-1\},$ (13)
where $g_{2}(r)$ is the pair-correlation function with the dimensionless distance $r$ normalized by the radius of each particle. If we adopt the Percus-Yevick approximation [53], the integral in Eq. (13) can be evaluated as $\int_{2}^{\infty}drr(g_{2}^{\mathrm{PY}}(r)-1)=-2(5-\Phi+\Phi^{2}/2)/5(1+2\Phi)$. Thus, we obtain $\langle M^{\infty}\rangle\approx(1-\Phi)^{3}/(1+2\Phi)$, which gives a reasonable sedimentation rate with the approximation [11]: $\langle M\rangle\approx(\langle M^{\infty}\rangle^{-1}+\langle R^{\mathrm{lub}}\rangle)^{-1}\approx(1-\Phi)^{3}/(1+2\Phi+1.492\Phi(1-\Phi)^{3})$ for $\Phi<0.49$. However, the calculation of $g_{2}(r)$ for $\Phi>0.49$ is difficult. Instead, we adopt the following empirical relation. It is known that $g_{2}(r)$ is zero for $r<2$, has a sharp peak at $r=2$, and oscillates around $1$ for $r>3$ for dense liquids. Thus, we assume the integral part is proportional to the radial distribution at contact as
$\int_{2}^{\infty}drr(g_{2}(r)-1)\approx\frac{g_{0}^{\mathrm{T}}(\Phi)}{g_{0}^{\mathrm{CS}}(\Phi)}\left\{\frac{\Phi(11-\Phi)}{5(1+2\Phi)}\right\},$ (14)
where $g_{0}^{\mathrm{T}}(\Phi)=(1-\Phi_{A}/2)(\Phi_{\mathrm{rcp}}-\Phi_{A})/\{(1-\Phi_{A})^{3}(\Phi_{\mathrm{rcp}}-\Phi\}$ introduced in Ref. [54] and $g_{0}^{\mathrm{CS}}(\Phi)=(1-\Phi/2)/(1-\Phi)^{3}$ introduced in Ref. [55] with $\Phi_{A}=0.49$ and $\Phi_{\mathrm{rcp}}=0.639$. Substituting Eq. (14) into Eq. (13), we obtain the empirical far-field contribution to the sedimentation rate for $\Phi>0.49$ as
$\langle M^{\infty}\rangle\approx 1+\Phi-\frac{\Phi^{2}}{5}+\frac{g_{0}^{\mathrm{T}}(\Phi)}{g_{0}^{\mathrm{CS}}(\Phi)}\left\{\frac{\Phi(11-\Phi)}{5(1+2\Phi)}\right\}.$ (15)
Then, the sedimentation rate can be obtained as
$\langle M\rangle\approx(\langle M^{\infty}\rangle^{-1}+\langle R^{\mathrm{lub}}\rangle)^{-1}.$ (16)
If we believe Eqs. (15) and (16), $\langle M\rangle\approx 2.8$ for $\Phi=0.52$ and $\langle M\rangle\approx 3.0$ for $\Phi=0.56$. This is an interesting result that supports the insensitivity of $\eta$ against $\Phi$.